On an old 45 rpm record, the beginning track is 8 cm from the center and the

end track is 5 cm from the center.

What are the angular speeds and

the tangential speeds at these distances when the record is spinning at 45 rpm?



1
Expert's answer
2022-03-07T18:06:28-0500

"\omega_1=\omega_2=2\pi\nu=4.71~\frac{rad}s,"

"v_1=\omega r_1=37.7~\frac{cm}s,"

"v_2=\omega r_2=23.6~\frac{cm}s."